# MultiBanana: A Challenging Benchmark for Multi-Reference Text-to-Image Generation

### v1

@@ E.3 Further Discussion of Difficult Reference Combination
Cross-domain diversity   In cross-domain cases, samples with reference images from different domains (darker color) were evaluated as lower Reference Consistency scores, because all models aim to unify the domains during generation.
Scale and viewpoint differences   In the different scale and viewpoint cases, the Reference Consistency scores declined consistently across all models. According to the reasoning traces of the VLM, the primary reasons for deductions are the loss of fine details and changes in pose. This suggests that when models attempt to reproduce objects under different scales or viewpoints, they may fail to preserve fine details or may adjust the pose to achieve a more natural appearance.
Rare concept references   In the rare concept case, because the reference images often contain subjects at a large scale, instances were observed in which the model reproduced this scale without adjustment, even when it was too large. For this reason, rare concept samples tend to receive lower scores on the Physical Realism.
Multilingual references   In the multilingual text rendering case, we observed a consistent decline across all metrics except Reference Consistency. This primarily occurs because models have limited text-rendering capability in languages other than English, leading to failures in cross-language conversion. As a result, instruction alignment and visual quality become lower. However, scores dropping in GPT-Image-1 were lower than those of others, since it exhibits comparatively stronger text-rendering ability in non-English languages.
E.4 Agentic Inference
@@
We introduced three agentic frameworks, Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA). Here, we refer to $\mathrm{Gen}$ as the Generator, which produces an image, and $\mathrm{Plan}$ as the Planner, which updates the instruction prompt (and selects reference images) for the next step. Let $P^{t}$ denote the instruction prompt at step $t$, $\{R_{i}\}_{i\in[I]}$ the reference images, and $G^{t}$ the generated image with $G^{0}=\varnothing$. We conducted experiments with both GPT and Gemini. For GPT, we use GPT-Image-1 as the Generator and GPT-5 as the Planner. For Gemini, we use Nano Banana as the Generator and Gemini-2.5-Flash as the Planner.
E.4.1 Iterative Prompt Refinement (IPR)
IPR framework is formulated as follows. The Planner refines the prompt based on the generation result from the previous step.
@@ E.4.1 Iterative Prompt Refinement (IPR)
E.4.2 Context-Aware Feedback Generation (CAFG)
CAFG framework is formulated as follows. The Generator generates the image based on the generation result (context) from the previous step.
@@ E.4.2 Context-Aware Feedback Generation (CAFG)
E.4.3 Selective Reference Adaptation (SRA)
SRA framework is formulated as follows. The Planner selects only the reference images that should be improved based on the generation result from the previous step, and the Generator receives these as context. SRA is expected to reduce task complexity for the agent by adaptively decreasing the number of references.
@@ E.4.3 Selective Reference Adaptation (SRA)
E.4.4 Experiments
We set the maximum number of steps $t$ as $3$ and we evaluate three agentic frameworks—IPR, CAFG, and SRA—using Gemini, and the IPR framework using GPT. Figure 22 illustrates an example of step-wise improvement in the IPR framework. Figure 23, Figure 24, Figure 25, and Figure 26 present the evaluation results of multi-reference generations at each step, averaged across judgments from Gemini-2.5-Flash and GPT-5, broken down by category. Figure 27 presents the results for single and two-reference generations. Nano Banan (Gemini) shows modest improvements in Physical Realism and Visual Quality across refinement steps, while Instruction Alignment and Reference Consistency either remain unchanged or deteriorate. In contrast, GPT demonstrates consistent improvements across all categories. This suggests that Gemini’s planner progressively loses information from the original prompt as refinement steps proceed.
@@ E.4.4 Experiments
Figure 27: Detailed results of single and two-reference image generation.
F Extended Related Works
F.1 Controllable Text-to-Image Generation
Diffusion models have achieved remarkable progress in generative modeling, demonstrating exceptional performance, particularly in high-fidelity image synthesis. Representative systems, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3], FLUX [flux2024, labs2025flux1kontext], DALL-E [ramesh2022dalle], and Imagen [saharia2022imagen, imagenteamgoogle2024imagen3], have demonstrated strong text-to-image generation capabilities, establishing a controllable and scalable foundation for image generation. To enhance controllability, models such as ControlNet [zhang2023controlnet] and T2I-Adapter [mou2023t2i-adapter] introduced external conditioning modules, enabling image-conditioned generation. Recent large pretrained diffusion models with textual conditioning, such as Stable Diffusion [rombach2022ldm, podell2024sdxl, esser2024sd3] and FLUX [flux2024, labs2025flux1kontext], enable the synthesis of novel scene images while preserving the identity depicted in the reference images. Building on these developments, approaches based on fine-tuning (e.g., DreamBooth [ruiz2022dreambooth]) and adapter-based methods (e.g., IP-Adapter [ye2023ip-adapter]) have been actively explored to address this challenge [jiang2025infiniteyou, gal2023an, brooks2023instructpix2pix, mokady2023null, tan2025ominicontrol, NEURIPS2023_602e1a5d]. Additionally, training-free approaches (e.g., pix2pix-zero [parmar2023zero] and prompt-to-prompt [hertz2023prompttoprompt]) have also been widely proposed [Cao_2023_ICCV, Huberman-Spiegelglas_2024_CVPR, Wallace_2023_CVPR, Tumanyan_2023_CVPR, kulikov2025flowedit, miyake2025negative-prompt-inversion]. These advancements collectively demonstrate that diffusion models are becoming increasingly flexible and adaptable across diverse conditional generation settings.
In recent years, research on unified image generation has gained momentum, aiming to handle multiple image generation tasks within a single model. For example, ChatGPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana] are capable of jointly processing text and image inputs, enabling integrated image generation and editing within a unified framework. Similarly, the open-source models, such as Qwen-Image [wu2025qwenimagetechnicalreport] and FLUX.1-Kontext-dev [labs2025flux1kontext], built upon a robust multimodal large language model backbone, demonstrate high-quality and flexible image generation and editing performance. Beyond these, numerous open-source unified generative models continue to emerge [xiao2024omnigen, xia2025dreamomni, wu2025omnigen2, xia2025dreamomni2].
With the advent of multimodal foundation models capable of jointly handling text and images [google2025nanobanana, openai2025gpt4oimage], a new task has recently become feasible: multi-reference-based image generation. In this task, the model receives multiple reference images along with textual instructions and generates a new scene while preserving the identity of the reference images. However, despite its novelty and potential, a reliable benchmark for evaluating this innovative task has yet to be established, leaving it as an essential open challenge for future research.
F.2 Benchmarks for Reference-Based Generation
Among reference-driven image generation tasks, the most widely recognized one is image editing. Image editing refers to the task of generating an edited image based on a single reference image and a textual instruction describing how the edit should be performed. Representative benchmarks such as MagicBrush [zhang2023magicbrush] and EMU-Edit [sheynin2024emuedit] have extended task-specific evaluations but still rely heavily on similarity-based metrics. SmartEdit [huang2024smartedit] supports editing of complex scenes, but does not sufficiently cover more general settings. Similarly, I2E-Bench [ma2024i2ebench] employs GPT-4o to provide human-aligned evaluations across a variety of editing tasks, yet it uses different metrics for each task and therefore fails to capture the shared characteristics of editing as a whole. More recently, ImgEdit [ye2025imgedit] has enabled the evaluation of multi-turn editing tasks, where image generation and editing are alternated interactively.
On the other hand, one of the earliest benchmarks for reference-based image generation beyond editing is the DreamBooth [ruiz2022dreambooth] dataset. However, such benchmarks are still limited to generation tasks conditioned on a single reference image and thus cannot handle more complex or compositional settings.
In recent years, following advances demonstrated by GPT-Image-1 [openai2025gpt4oimage] and Nano Banana [google2025nanobanana], increasing attention has been directed toward multi-reference image generation, where multiple reference images are used jointly for generation. Benchmarks such as OmniContext [wu2025omnigen2] and DreamOmni2 [xia2025dreamomni2] have been proposed to evaluate this task. Nevertheless, these benchmarks encompass only a small number of images and task types, thereby limiting the scope of the evaluation. Moreover, despite being designed for multi-reference generation, they do not evaluate qualitative differences or relationships between reference images, and their setups remain similar to those of conventional single-image editing benchmarks.
G Prompts
G.1 Prompts of Agentic Framework
G.2 Prompts of VLM as Judge
[…]

### v2

@@ F.3 Difficult Reference Combination
Cross-domain diversity   In cross-domain cases, samples with reference images from different domains (darker color) received lower Reference Consistency scores because all models tend to distort individual reference attributes to achieve a consistent style.
Scale and viewpoint differences   In the different scale and viewpoint cases, the Reference Consistency scores declined consistently across all models. This suggests that when models attempt to reproduce objects at different scales or from different viewpoints, they may fail to preserve fine details or adjust the pose to achieve a more natural appearance.
Rare concept references   In the rare-concept case, the model may struggle to handle uncommon subjects relative to more familiar ones. Because the reference images often depict these subjects at a large scale, the model tends to reproduce this scale without appropriate adjustment. This suggests that, unlike common concepts, the model may find it more difficult to flexibly control attributes such as size when dealing with rare concepts. As a result, rare concept samples tend to receive lower Physical Realism scores.
Multilingual references   In the multilingual text rendering case, we observed a consistent decline across all metrics except Reference Consistency. This may occur because models have limited text-rendering capability in languages other than English, leading to failures in cross-language conversion. As a result, instruction alignment and visual quality decrease. However, GPT-Image-1 exhibits a smaller decline than other models, owing to its comparatively stronger text-rendering ability in non-English languages.
F.4 Risks of Model Bias in Synthetic References
Synthetic references generated by closed models (i.e., Nano Banana, GPT-Image-1) may introduce in-distribution advantages when those same models are also evaluated on the benchmark. To verify this, we divided the benchmark into three subsets based on the source of the reference images: those from GPT-Image-1, those from Nano Banana, and those from both. The mean scores across all subsets remain within their respective 99% confidence intervals, indicating no statistically significant bias (Figure 12).
Figure 12: Analysis of potential in-distribution bias. Error bars indicate the 99% confidence intervals. For image generation models, we use the versions available as of January 2026.
F.5 Potential Conflict in Cross-Domain Task
One might expect that preserving subject details (reference consistency) and maintaining a coherent scene (background-subject match) are inherently at odds in cross-domain tasks—e.g., placing a photorealistic person into an anime background. However, our evaluation shows this is not always the case. As shown in Figure 13 (left), a generated image can satisfy both criteria, achieving a GPT-5 score of 7 for reference consistency and 9 for background-subject match. To further verify this, we created cross-domain editing tasks with explicit instructions that prioritize one criterion over the other for 20 randomly selected prompts. Table 15 shows that Nano Banana can improve the prioritized criterion while maintaining the other criterion.
Table 15: Comparison of the mean scores when prioritizing either reference consistency or background–subject match for the subset generated by Nano Banana, latest version as of January 2026.
Prompt
Reference
Consistency
Background–
Subject Match
Original 3.58 3.67
Consistency priority 4.17 3.25
Matching priority 3.40 4.03
Figure 13: Left: A sample generated by GPT-Image-1 on a cross-domain task. Right: Task example of our benchmark directly relevant to real-world use cases, such as advertising.
F.6 Alignment with Real-World Use Cases
As shown in Figure 13 (right), MultiBanana is directly relevant to real-world use cases, such as advertising. Our benchmark development using synthetic data also aligns with the community standard. For example, ImgEdit [91] similarly employs GPT-4o for instruction generation, which also aims to develop “practical and powerful tools for real-world applications.”
F.7 Agentic Inference
@@
We introduced three agentic frameworks: Iterative Prompt Refinement (IPR), Context-Aware Feedback Generation (CAFG), and Selective Reference Adaptation (SRA). Here, we refer to $\mathrm{Gen}$ as the Generator, which produces an image, and $\mathrm{Plan}$ as the Planner, which updates the instruction prompt (and selects reference images) for the next step. Let $P^{t}$ denote the instruction prompt at step $t$, $\{R_{i}\}_{i\in[I]}$ the reference images, and $G^{t}$ the generated image with $G^{0}=\varnothing$. We conducted experiments with both GPT and Gemini. For GPT, we use GPT-Image-1 as the Generator and GPT-5 as the Planner. For Gemini, we use Nano Banana as the Generator and Gemini 2.5 Flash as the Planner.
F.7.1 Iterative Prompt Refinement (IPR)
The IPR framework is formulated as follows. The Planner refines the prompt based on the generation result from the previous step.
@@ F.7.1 Iterative Prompt Refinement (IPR)
F.7.2 Context-Aware Feedback Generation (CAFG)
The CAFG framework is formulated as follows. The Generator generates the image based on the generation result (context) from the previous step.
@@ F.7.2 Context-Aware Feedback Generation (CAFG)
F.7.3 Selective Reference Adaptation (SRA)
The SRA framework is formulated as follows. The Planner selects only the reference images that should be improved based on the generation result from the previous step, and the Generator receives these as context. SRA is expected to reduce agents’ task complexity by adaptively decreasing the number of references.
@@ F.7.3 Selective Reference Adaptation (SRA)
F.7.4 Experiments
We set the maximum number of steps $t$ to $3$ and evaluate three agentic frameworks—IPR, CAFG, and SRA—using Gemini, and the IPR framework using GPT. Figure 14 illustrates an example of step-wise improvement in the IPR framework. Figure 20, Figure 21, Figure 22, and Figure 23 present the evaluation results of multi-reference generations at each step, averaged across judgments from Gemini 2.5 Flash and GPT-5, broken down by category. Figure 24 presents the results for single and two-reference generations. Nano Banana (Gemini) shows modest improvements in Physical Realism and Visual Quality across refinement steps, while Instruction Alignment and Reference Consistency either remain unchanged or deteriorate. In contrast, GPT demonstrates consistent improvements across all categories. This suggests that Gemini’s planner progressively loses information from the original prompt as refinement steps proceed.
G Implementations
API endpoints: For VLM evaluation, we used stable API versions: gpt-5-2025-08-07 and gemini-2.5-flash. For image generation models, we use the latest available API versions as of November 2025, unless otherwise noted.
H Extended Related Works
H.1 Controllable Text-to-Image Generation
Stable Diffusion [65, 60, 15], FLUX [42, 41], DALL-E [63], and Imagen [67, 32], have demonstrated strong text-to-image generation capabilities, establishing a scalable foundation for the task. To enhance controllability, models such as ControlNet [94] and T2I-Adapter [52] introduced external conditioning modules, enabling image-conditioned generation. Additionally, training-free approaches (e.g., pix2pix-zero [59] and prompt-to-prompt [25]) have also been widely proposed [6, 31, 75, 74, 39, 49]. These advancements collectively demonstrate that diffusion models are becoming increasingly flexible and adaptable across diverse conditional generation settings.
H.2 Benchmarks for Reference-Based Generation
Reference-based image generation is closely related to industrial applications such as content production [66, 90, 78, 96, 88, 37], advertising [33, 28, 51], and fashion design [98, 11, 16, 16, 12, 86, 36]. Among reference-driven image generation tasks, the most widely recognized one is image editing. RealEdit [71] focuses on single-image editing and provides an empirical analysis of real-world image-editing use cases. SpotEdit [20] proposes a benchmark for visually guided image editing. These works focus on benchmarking single-reference image edition/generation, while we focus on multi-reference image generation. In recent years, increasing attention has been directed toward multi-reference image generation, in which multiple reference images are jointly used to generate a single image. MultiRef-Bench [9] studies controllable image generation with multiple visual references, but our benchmark outperforms it by providing 3,769 samples (compared to 1,990), supporting up to 8 references (compared to 6), and using raw images (compared to bounding boxes or masks), making it more practical.
H.3 Instruction-Following in LLMs
The task of generating an image that faithfully reflects multiple references is analogous to instruction following in LLMs [97, 95, 46, 14, 76], in which models must simultaneously satisfy multiple constraints. In the text domain, recent studies have revealed that LLMs struggle to follow all given instructions as the number of constraints grows, often ignoring or inadequately addressing a subset of them [35, 24, 23, 40]. Multi-reference image generation poses a parallel challenge: as the number of reference images increases, models must reconcile potentially conflicting visual cues while remaining faithful to every reference.
I Limitations and Discussions
Our benchmark focuses on static image generation and does not address video generation or editing [27, 21, 55], where temporal consistency across frames introduces additional challenges for reference-based conditioning [29, 10, 1]. Moreover, our prompts describe object placement in natural language (e.g., ”on the left,” ”in the foreground”), which is inherently ambiguous. Combining reference images with structured layout inputs [94, 89, 45, 61] could enable finer-grained control over subject composition.
J Prompts
J.1 Prompts of Agentic Framework
J.2 Prompts of VLM as Judge
[…]
Figure 17: Average scores of each model across evaluation metrics for each task by the average of GPT and Gemini. The horizontal axis denotes the scores for the five evaluation criteria, and the vertical axis denotes the number of reference images.
Figure 18: Total scores of each model for single and two-reference tasks by GPT, Gemini, and their average.
Figure 19: Total scores for each difficult reference combination across models. Darker colors represent the average score for tasks that include the corresponding combination, whereas lighter colors indicate the average score for tasks that do not include it.
@@ I Limitations and Discussions
Figure 24: Detailed results of single and two-reference image generation using the agentic frameworks.
Figure 25: Qualitative example for 4-Object Tasks.
Figure 26: Qualitative example for 3-Object + Background Tasks.
Figure 27: Qualitative example for 3-Object + Local Tasks.
Figure 28: Qualitative example for 3-Object + Global Tasks.
Figure 29: Qualitative example for Tasks with 8 references.